OpenStudy (anonymous):

Are two rational expressions equivalent if they have the same simplified form?

OpenStudy (calculusxy):

Yes if you say that you are working with two ratios of 5:10 and 10:20, when you simplify fully you will get: 5/10=1/2 and 10/20=5/10=1/2

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank you guys so much @calculusxy